OpenStudy (anonymous):

Sunset rents an SUV at $21.95 plus $0.23 per mile. Sunrise rents the same vehicle for $24.95 plus $0.19 per mile. For what mileage is the cost the same?

OpenStudy (anonymous):

help plzz

OpenStudy (anonymous):

21.95 + .23m = 24.95 + .19m now solve for m and that will be your answer

OpenStudy (anonymous):

is it 42

OpenStudy (anonymous):

then i got 46.90 then do i divivde

OpenStudy (anonymous):

.23m -.19m = 24.95 - 21.95 .04m = 3 m = 3/.04 m = 75

OpenStudy (anonymous):

can you help me with more plzzz and is that the finial answer

OpenStudy (anonymous):

the final answer is 75 . What other questions do you have ?

OpenStudy (anonymous):

Six pears and three apples cost $3.90. Two pears and five apples cost $3.30. How much does one pear cost?

OpenStudy (anonymous):

I am sorry...I will be right back in about 30 minutes....I have to go somewhere

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Pears are $.40 and apples are $.50

OpenStudy (anonymous):

A kayaker paddled 2 hours with a 6 mph current in a river. The return trip against the same current took 3 hours. Find the speed the kayaker would make in still water.

OpenStudy (anonymous):

and thank you

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

Your piggy bank has 25 coins in it; some are quarters and some are nickels. You have $3.45. How many nickels do you have? (I know you can figure this out by trial and error, but you must write the equations and solve.)

OpenStudy (anonymous):

There are 11 quarters and 14 nickels in the piggy bank lolxx u got the easy part of algebra2

OpenStudy (anonymous):

lol i got more

OpenStudy (anonymous):

There were 1500 people at a high school football game. Student tickets were $2.00 and adult tickets were $3.50. The total receipts for the game were $3825. How many students bought tickets?

OpenStudy (anonymous):

An airplane flew 4 hours with a 25 mph tail wind. The return trip against the same wind took 5 hours. Find the speed of the airplane in still air.

OpenStudy (anonymous):

s + a = 1500 -->a = 1500 - s 2s + 3.5a = 3825 now sub 1500 - s in for a in the 2nd equation 2s + 3.5(1500 - s) = 3825 2s + 5250 - 3.5s = 3825 - 3.5s + 2s = 3825 - 5250 -1.5s = -1425 s = -1425/-1.5 s = 950 student tickets were bought

OpenStudy (anonymous):

i got the answer for that already lol sorry can you help me with the airplane one

OpenStudy (anonymous):

what a coincidence :)

OpenStudy (anonymous):

omg so the answer is225 lol

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

too bad we can't find more of your questions that way...lol

OpenStudy (anonymous):

A kayaker paddled 2 hours with a 6 mph current in a river. The return trip against the same current took 3 hours. Find the speed the kayaker would make in still water.

OpenStudy (anonymous):

how bout this one

OpenStudy (anonymous):

????

OpenStudy (anonymous):

s + 6 (with the current) s - 6 (against the current) 2(s + 6) = 3(s - 6) 2s + 12 = 3s - 18 2s - 3s = -18 + 12 -s = - 6 s = 6 mph.....I think